produced Calculate K. AgCl is shaken with water at 25°C, a solution containing 1.0 X 10M silver ions is
Answer #1

Ksp (Agce) = [agti [ce] Agce. - Agt + ce- [Ag+] = [ck] = 1.0X10-5 m Ksp = 10-5 x 10-5 m2 = 2x ixio-10 m2 Canced amscaner